I am having a terrible time solving this equation 4^2 - 4x = -3
It must be solved using the Indian method of first multiplying each term by 4. I have done that with the result of 16^2 - 16x = -12. The next step in this method is to add the square of the coefficient of the original x term to both sides, which results in 16^2 - 16x + 16 = -12 + 16. I have this correct, but cannot figure out how to complete the equation. How can I factor this equation, and then solve it.
I am pulling my hair out!
Ellen

Answer by mangopeeler07(462)

16^2 - 16x + 16 = -12 + 16

Factor out a 16 on the left side
16(16-x+1)=-12+16

Simplify the parentheses
16(17-x)=-12+16

Distribute
272-16x=-12+16

Combine like terms
272-16x=4

Subtract 272 from both sides
-16x=-268

Divide by -16
x=16.75